Suppose supply of a good is perfectly elastic at a price of $5. The market demand curve for this good is linear, with zero quantity demanded at a price of $25. Given that the slope of this linear demand curve is -0.25, draw a supply and demand graph to illustrate the consumer surplus that occurs when the market is in equilibrium.
1
Expert's answer
2014-06-11T08:29:12-0400
If Ed = -0.25 and quantity demanded is zero at P = $25, then Qd = 25 -0.25P.

As Qs = 5, in equilibrium Pe = $5. Qe = 25 - 0.25*5 = 23.75 units.

Consumer surplus = 0.5*23.75*(25 - 5) = $237.5.